[The following information applies to the questions displayed below.]

Warner Clothing is considering the introduction of a new baseball cap for sales by local vendors. The company has collected the following price and cost characteristics.

Sales price $ 18 per unit
Variable costs 2 per unit
Fixed costs 52,000 per month

a. What number must Warner sell per month to break even?

b. What number must Warner sell per month to make an operating profit of $40,000?


Break-even sales in units b. Number of units to be sold

Assume that the company plans to sell 7,000 units per month. Consider requirements (b), (c), and (d) independently of each other.

Required:

a. What will be the operating profit?

b. What is the impact on operating profit if the sales price decreases by 10 percent? Increases by 20 percent?

c. What is the impact on operating profit if variable costs per unit decrease by 10 percent? Increase by 20 percent?

d. Suppose that fixed costs for the year are 10 percent lower than projected, and variable costs per unit are 10 percent higher than projected. What impact will these cost changes have on operating profit for the year? Will profit go up? Down? By how much?

Required A Required B Required C Required D What will be the operating profit? Operating profitRequired A Required B Required c Required D What is the impact on operating profit if the sales price decreases by 10 percentRequired A Required B Required C Required D What is the impact on operating profit if variable costs per unit decrease by 10Required A Required B Required C Required D Suppose that fixed costs for the year are 10 percent lower than projected, and va
Answer #1

a. What number must Warner sell per month to break even?

Answer : Break-even sales in unit : 3,250

Calculation:

Profit = (P – V)X – F

$0 = ($18 – $2)X – $52,000

$16X = $52,000

X =$52,000/$16

X = 3,250 units

b. What number must Warner sell per month to make an operating profit of $40,000?

Answer : Number of units to be sold: 5,750

Calculation:

Profit = (P – V)X – F

$40,000 = ($18 – $2)X – $52,000

$16X = $92,000

X =$92,000/$16

X = 5,750 units

Assume that the company plans to sell 7,000 units per month. Consider requirements (b), (c), and (d) independently of each other.

Required:

a. What will be the operating profit?

Answer : Operating Profit: 60,000

Calculation:

Profit = ($18 – $2) × 7,000 – $52,000

= $60,000.

b. What is the impact on operating profit if the sales price decreases by 10 percent? Increases by 20percent?

Answer : sales price decreases by 10 percent,  Operating Profit decreases by $12,600

sales price Increases by 20percent Operating Profit increases by $25,200

Calculation:

10% price decrease.

Now P = $16.20

Profit = ($16.20 – $2) × 7,000 – $52,000 = $47,400.

Profit decreases by $12,600

20% price increase.

Now P = $21.60

Profit = ($21.60 – $2) × 7,000 – $52,000 = $85,200.

Profit increases by $25,200

c.What is the impact on operating profit if variable costs per unit decrease by 10 percent? Increase by 20 percent?

Answer : variable costs per unit decrease by 10 percent,  Operating Profit increases by $1,400

variable costs per unit Increases by 20 percent Operating Profit decreases by $2,800

Calculation:

10% variable cost decrease.

Now V = $1.80

Profit = ($18 – $1.80) × 7,000 – $52,000 = $61,400.

Profit increases by $1,400

20% variable cost increase.

Now V = $2.40

Profit = ($18 – $2.40) × 7,000 – $52,000 = $57,200.

Profit decreases by $2,800

d.Suppose that fixed costs for the year are 10 percent lower than projected, and variable costs per unit are 10 percent higher than projected. What impact will these cost changes have on operating profit for the year? Will profit go up? Down? By how much?

Answer : Operating Profit increases by $3,800

Calculation:

Profit = ($18 – $2.20) × 7,000 – $46,800= $63,800.

Profit increases by $3,800